OpenStudy (anonymous):

Simplify the following quotient of complex numbers into the form a + bi. 1+8i/-6+i

OpenStudy (anonymous):

1-i/3

OpenStudy (angela210793):

multiply up and down with (-6-i)---> (1+8i)*(-6-i)/35=(2-49i)/35

OpenStudy (anonymous):

2/37-(49 i)/37 I believe twoof is wrong

OpenStudy (anonymous):

okay

OpenStudy (angela210793):

pk u have 36-1 not 36+1

OpenStudy (zarkon):

pk51 is correct

OpenStudy (angela210793):

don't u have (6-1)*(6+1)=6^2-1^2=36-1=35? :S

OpenStudy (zarkon):

no

OpenStudy (anonymous):

oops, I solved it like there were no parenthesis

OpenStudy (angela210793):

can u explain me pls?

OpenStudy (zarkon):

\[(-6+i)(-6-i)=(-6^2)-i^2=36-(-1)=37\]

OpenStudy (angela210793):

righttt Thnx ^_^